A well celebrated product sold 13 million units as of June 2017. The product retails for $249. Suppose Ali is willing to pay $280, Mona is willing to pay $225, and Bodie is willing to pay $350 for the product. What is the value of the market consumer surplus?

A)

$132

B)

$855

C)

$630

D)

$498

E)

$250
Answer #1

Answer
Individual consumer surplus =williness to pay -price
Price=$249

Ali's CS =280-249
=$31

Mona's willingness to pay is lower than the price so she can't buy it or can not afford to buy it she is not considered in the market.

Bodie's consumer surplus =350-249
=$101
total surplus =sum of individual surpluses
=31+101
=132
the total consumer surplus is $132

option A